# Identifying treatment response subgroups in observational time-to-event data

### v3

@@ 5 Conclusion
This work fills the current gap in the literature of methodologies to identify subgroups of treatment responses in observational time-to-event data. While modelling the observed outcome, the approach jointly identifies patient subgroups with different treatment responses. Our experiments demonstrate the capacity of our proposed method to uncover groups of treatment effects in both synthetic and real-world datasets. However, causal modelling relies on empirically unverifiable assumptions. The invalidity of these assumptions is a lesser concern in our targeted setting of hypothesis generation, in comparison to individual estimation of treatment recommendations.
Serving the purpose of a hypothesis-generating tool, we invite practitioners to further investigate the subgroups identified in observational studies through RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
@@ Acknowledgements
The authors acknowledge the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2). VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1. FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z).
@@ A.1 Individualised Treatment Effect
This section derives the individualised treatment effect expression introduced in equation 3.
@@
$=\mathbb{E}(\mathbbm{1}(T_{1}\geq t)\mid X=x)-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
$=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under assumption 4 and 5)
A.2 Upper-Bound Simplification
In Section 2, we claim that under the considered DAG presented in Figure 2,
$\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$
with $A$, being the treatment assignment. Under assumption 3, expressing this quality corresponds to:
$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$
with $q$, the density function. From this expression, $q_{Z\mid X}^{A=0}=q_{Z\mid X}^{A=1}$, which results in the distance between these distributions being null.
@@ B.1 Data generation
We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to Nagpal et al. (2022a). The following describes our generation process:
@@
$T_{0}\mid Z,X,B^{0}_{z},\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
@@
$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{1}_{z}=(z,x,\beta_{z}^{1},\gamma_{z}^{1})$ $\sim\text{Gompertz}\left(w_{1}(\beta_{z}^{1},x),s_{1}(\gamma_{z}^{1},x)\right)$
with $w_{0}$, $w_{1}$ two functions parametrising the Gompertz distributions’ shape as $w_{0}(\beta,x):=|\beta[0]|+(x[5:10]\cdot\beta[5:10])^{2}$, $w_{1}(\beta,x):=|\beta[0]|+(x[1:5]\cdot\beta[1:5])^{2}$, and the shift parameter parameterised as $s_{0}(\gamma,x):=|\gamma[0]|+|(x[1:5]\cdot\gamma[1:5])|$ and $s_{1}(\gamma,x):=|\gamma[0]|+|(x[5:10]\cdot\gamma[5:10])|$ where $v[i]$ described the $i^{th}$ element of the vector $v$. These functions aim to introduce non-linear responses with discrepancies between control and treatment regimes. Note that we allow covariates to influence the survival distribution as a patient’s covariates influence Gompertz’s shapes and scales.
@@
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world applications. Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as "Randomised" and "Observational". "Randomised" consists of a Bernoulli draw using the realisation of $P$. "Observational" reflects an assignment dependent upon the observed covariates.
$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(0.5)$
$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times 0.5)$
with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$. In our experiment, we chose $\Phi(x)=\sum x^{2}$.
@@
From these data, the goal is to model the treatment effect from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)*t_{0}+a*t_{1})$ and $D=\mathbbm{1}_{C>T}$.
@@ B.2 Training and hyperparameter optimisation
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2. All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
@@
We adopted a 100-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[10,25,50]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $100$ or $250$.
B.3 Sensitivity
In Section 3, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$. This section explores how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters. This analysis examines the risk associated with misspecifying the number of clusters.
Figure 5 illustrates the cross-validated subgroup treatment effects when trained with $K=2,3$ and $4$ clusters. Uncertainty increases with misspecified models due to the instability of the identified subgroups.
(a) $K=2$
(b) $K=3$ (as in the main text)
(c) $K=4$
B.4 Alternative data generations
In this section, we explore alternative data generations to demonstrate the robustness of the proposed strategy.
B.4.1 Unequal cluster size
Section B.1 presents a population equally distributed over the different clusters. In medical settings, the underlying subgroups may differ in size. This section explores an alternative scenario in which the population is distributed over the three clusters: 62.5%, 25% and 12.5%.
Table 2 summarises the performance in this simulation, echoing the main text’s conclusions. The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects..
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.637 (0.068) 0.064 (0.013) 0.021 (0.009) 0.062 (0.010) CSC Unadjusted 0.651 (0.049) 0.055 (0.009) 0.017 (0.010) 0.067 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.780 (0.086) 0.064 (0.027) 0.040 (0.011) 0.069 (0.010) CSC Unadjusted 0.616 (0.146) 0.072 (0.020) 0.039 (0.028) 0.117 (0.033) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
@@ B.4.1 Unequal cluster size
B.4.2 Number of points
Section B.1 describes a data generation with $3,000$ patients. This section presents results when there are $300$ and $30,000$ patients. A smaller number of patients may result in a larger impact of treatment non-randomisation but also impact the neural network capacity to identify underlying subgroups of treatment effect. A larger number makes non-randomisation less of a concern.
@@ B.4.2 Number of points
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.504 (0.139) 0.131 (0.090) 0.085 (0.033) 0.091 (0.062) CSC Unadjusted 0.449 (0.106) 0.202 (0.049) 0.162 (0.055) 0.125 (0.088) CMHE ($K=3$) 0.126 (0.084) 0.340 (0.050) 0.135 (0.043) 0.166 (0.050) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.087 (0.118) 0.284 (0.056) 0.639 (0.939) 0.747 (1.014) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.774 (0.015) 0.018 (0.010) 0.012 (0.003) 0.030 (0.009) CSC Unadjusted 0.737 (0.022) 0.033 (0.006) 0.010 (0.002) 0.039 (0.008) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
B.4.3 Number of clusters
Section B.1 describes a data generation with $K=3$ clusters. This section explores the behaviour of the different methodologies with an increased number of clusters $K=5$. Following the same data generation, we introduces two additional clusters with centres [-3, 3] and [4, 4].
Table 4 presents the different models’ performance in this setting, echoing the same results as the main text. The proposed method best recovers the underlying clustering structure as shown by the Rand-Index. CMHE suffers from its inherent association between treatment and survival. The step-wise approach is unable to identify the clusters of interest if the data presents a structure that is not fully aligned with the outcome of interest.
This scenario further validates our findings, demonstrating the methodology’s capacity to generalise beyond the scenario presented in the main text.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ Randomised CSC 0.475 (0.038) 0.073 (0.030) 0.059 (0.023) 0.041 (0.015) 0.099 (0.024) 0.017 (0.003) CSC Unadjusted 0.459 (0.037) 0.095 (0.055) 0.054 (0.027) 0.046 (0.011) 0.076 (0.020) 0.019 (0.005) CMHE ($K=5$) 0.306 (0.016) 0.091 (0.036) 0.104 (0.012) 0.015 (0.002) 0.093 (0.011) 0.232 (0.013) CMHE ($L=5$) 0.256 (0.018) 0.135 (0.020) 0.112 (0.008) 0.020 (0.003) 0.221 (0.009) 0.349 (0.006) CMHE ($K=L$) 0.372 (0.018) 0.101 (0.023) 0.181 (0.013) 0.072 (0.003) 0.119 (0.010) 0.323 (0.010) KMeans + TE 0.156 (0.012) 0.075 (0.009) 0.184 (0.016) 0.072 (0.012) 0.147 (0.014) 0.039 (0.010) Virtual Twins 0.284 (0.081) 0.034 (0.029) 0.066 (0.026) 0.029 (0.012) 0.082 (0.023) 0.036 (0.010) Observational CSC 0.446 (0.016) 0.119 (0.072) 0.087 (0.018) 0.026 (0.008) 0.097 (0.023) 0.027 (0.016) CSC Unadjusted 0.444 (0.048) 0.077 (0.046) 0.065 (0.026) 0.045 (0.020) 0.118 (0.026) 0.034 (0.022) CMHE ($K=5$) 0.303 (0.010) 0.095 (0.030) 0.112 (0.017) 0.018 (0.005) 0.094 (0.013) 0.229 (0.010) CMHE ($L=5$) 0.217 (0.026) 1.341 (2.717) 0.119 (0.014) 0.022 (0.005) 1.432 (2.720) 1.561 (2.709) CMHE ($K=L$) 0.371 (0.029) 0.105 (0.009) 0.208 (0.043) 0.076 (0.006) 0.186 (0.025) 0.325 (0.004) KMeans + TE 0.162 (0.011) 0.058 (0.005) 0.210 (0.031) 0.084 (0.018) 0.122 (0.032) 0.032 (0.006) Virtual Twins 0.320 (0.084) 0.062 (0.038) 0.026 (0.008) 0.044 (0.021) 0.077 (0.040) 0.038 (0.018)
Table 4: Cross-validated performance (with standard deviation in parenthesis) when $K=5$.
B.4.4 Impact of treatment rate

### v4

@@ 5 Conclusion
Understanding heterogeneity in treatment effect is essential for guiding clinical practice and informing the development of new treatments. Motivated principally by how real-world clinical guidelines depend on patient subgroups, this work introduces a neural network-based framework for identifying treatment-response subgroups in observational time-to-event data. By leveraging large-scale observational datasets and survival analysis with causal inference methods, our approach addresses a key gap in the literature: it moves beyond average or individualised treatment-effect estimation to reveal subgroups of treatment response.
Our empirical results on both synthetic and real-world datasets demonstrate that the proposed method successfully characterises latent subgroups with divergent treatment effects. While we acknowledge that our model depends on empirically unverifiable assumptions as often in causal inference, our focus on subgroup detection for hypothesis generation lessens this concern. Indeed, such subgroup identification serves as a starting point for more targeted RCTs, where further validation can (i) confirm the heterogeneity in responses, (ii) explore alternative treatment strategies in less responsive subgroups, and (iii) shape evidence-based clinical guidelines.
In doing so, this work contributes a flexible and practical tool for the design of future RCTs. By informing both the selection of patient populations most likely to benefit and avoiding treatments with limited or harmful effects, our framework offers a meaningful hypothesis-generating tool for clinicians, policymakers, and researchers striving to optimise patient care and resource allocation.
@@ Acknowledgements
The authors acknowledge the partial support of the UKRI Medical Research Council (programme numbers MC_UU_00002/5 and MC_UU_00002/2 and theme number MC_UU_00040/02 – Precision Medicine). VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1. FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z).
@@ A.1 Individualised Treatment Effect
This section derives the individualised treatment effect expression introduced in Section 3.
@@
$=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)\mid X=x)$
$-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
$=\mathbb{P}($ $T^{\prime}\geq t\mid A=1,X=x)$
$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under assumption 1 and 3)
@@ B.1 Data generation
We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to (Nagpal et al., 2022a). The following describes our generation process:
@@
$T_{0}\mid Z,X,B^{0}_{z},$ $\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$
$\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
@@
$T_{1}\mid Z,X,B^{1}_{z},$ $\Gamma^{1}_{z}=(z,x,\beta_{z}^{1},\gamma_{z}^{1})$
$\sim\text{Gompertz}\left(w_{1}(\beta_{z}^{1},x),s_{1}(\gamma_{z}^{1},x)\right)$
with $w_{0}$, $w_{1}$ two functions parametrising the Gompertz distributions’ shape as $w_{0}(\beta,x):=|\beta[0]|+(x[5:10]\cdot\beta[5:10])^{2}$, $w_{1}(\beta,x):=|\beta[0]|+(x[1:5]\cdot\beta[1:5])^{2}$, and the shift parameter parametrised as $s_{0}(\gamma,x):=|\gamma[0]|+|(x[1:5]\cdot\gamma[1:5])|$ and $s_{1}(\gamma,x):=|\gamma[0]|+|(x[5:10]\cdot\gamma[5:10])|$ where $v[i]$ described the $i^{th}$ element of the vector $v$. These functions aim to introduce non-linear responses with discrepancies between control and treatment regimes. Note that we allow covariates to influence the survival distribution as a patient’s covariates influence Gompertz’s shapes and scales.
@@
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world studies. Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as Randomised and Observational. Randomised consists of a Bernoulli draw using the realisation of $P$. Observational reflects an assignment dependent upon the observed covariates.
$A_{rand}$ $\sim\text{Bernoulli}(0.5)$
$A_{obs}\mid X=x$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times 0.5)$
with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$. In our experiment, we chose $\Phi(x)=\sum x^{2}$ for a non-linear treatment assignment.
@@
Our goal is to model the treatment effect and identify the underlying subgroup structure $Z$ from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)\times t_{0}+a\times t_{1})$ and $D=\mathbbm{1}_{C>T}$.
@@ B.2 Training and hyperparameter optimisation
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search. All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
@@
We adopted a 500-iteration random grid-search over the following hyperparameters: network depth between 0 and 3 inner layers with 50 nodes, latent subgroup representation in $[10,25,50]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $100$ or $250$.
The same set of parameters, when appropriate, was used for CMHE. Tree-based approach to estimate treatment effect had a depth limited to: 3, 5 or unconstrained, and the minimum sample split was chosen between 6, 12 or 24.
B.3 Sensitivity to choice of K
In Section 4, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$. While this outcome-guided selection of the number of clusters is a core strength of the proposed methodology in comparison to existing strategies, we explore how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters. This analysis examines the risk associated with misspecifying the number of clusters.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Obs. CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CSC Linear 0.819 (0.041) 0.047 (0.015) 0.031 (0.022) 0.022 (0.003)
Table 2: Cross-validated performance (with standard deviation in parenthesis).
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.643 (0.067) 0.060 (0.016) 0.021 (0.007) 0.070 (0.013) CSC Unadjusted 0.643 (0.071) 0.064 (0.013) 0.018 (0.008) 0.060 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.788 (0.113) 0.040 (0.017) 0.041 (0.017) 0.086 (0.043) CSC Unadjusted 0.626 (0.170) 0.080 (0.015) 0.031 (0.009) 0.095 (0.049) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
@@ B.3 Sensitivity to choice of K
Figure 5 illustrates the cross-validated subgroup treatment effects when the model is trained with $K=2,3$ and $4$ clusters, despite the existence of 3 clusters. When using a number of clusters different than indicated by the elbow heuristic, the identified clusters are less stable as not capturing the true underlying treatment responses, as shown by the increased standard deviation. Uncertainty increases with misspecified models, providing an additional tool to select the appropriate number of cluster: the stability of the estimated clusters over cross-validation indicates a better alignment with underlying distribution.
B.4 Linear Modelling
While the flexibility of neural network allows flexibility of the survival distribution, one could consider more interpretable assignment $G$ and treatment $W$ components than the neural network approach proposed in the main text. This section explores the recovery of treatment effects and subgroups when using linear regressions, i.e. no hidden layers in a neural network, instead of the previous neural networks.
Table 2 summarises the results presented in the main paper with the additional linear alternative in which both the treatment and assignment networks are replaced by a linear model, referred as CSC Linear. These results demonstrate that enforcing linear relation between covariates and the different quantities of interest can reduce the risk of overfitting in small datasets as shown by the improved Rand-Index. Note that under larger dataset or more complex relation between covariates, group membership and treatment, one should consider more flexible modelling, such as the proposed neural approach.
B.5 Alternative data generations
In this section, we explore alternative data-generative processes to demonstrate the robustness of the proposed strategy under different settings.
B.5.1 Unequal cluster size
Section B.1 presents an equally-distributed population over the different clusters. In medical settings, the underlying subgroups may differ in size. This section explores an alternative scenario in which the population is distributed over the three clusters as follows: 62.5%, 25% and 12.5%.
Table 3 summarises the performance in this simulation, echoing the main text’s conclusions. The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.310 (0.187) 0.236 (0.103) 0.110 (0.033) 0.149 (0.106) CSC Unadjusted 0.443 (0.285) 0.176 (0.098) 0.087 (0.046) 0.132 (0.112) CMHE ($K=3$) 0.113 (0.100) 0.296 (0.040) 0.137 (0.042) 0.188 (0.080) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.079 (0.119) 0.359 (0.068) 0.182 (0.089) 0.240 (0.058) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.764 (0.015) 0.027 (0.008) 0.011 (0.001) 0.047 (0.008) CSC Unadjusted 0.729 (0.022) 0.033 (0.008) 0.010 (0.003) 0.039 (0.009) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.830 (0.047) 0.051 (0.008) 0.023 (0.013) 0.020 (0.005) CSC Unadjusted 0.857 (0.033) 0.045 (0.006) 0.030 (0.015) 0.015 (0.003) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.134 (0.090) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.718 (0.198) 0.061 (0.050) 0.057 (0.048) 0.023 (0.005) CSC Unadjusted 0.804 (0.045) 0.037 (0.006) 0.023 (0.008) 0.025 (0.007) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.318 (0.140) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012)
Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates under observational settings.
B.5.2 Number of points
Section B.1 describes a data generation with $3,000$ patients. This section presents results when there are $300$ and $30,000$ patients. A smaller number of patients may impact the neural network capacity to identify underlying subgroups of treatment effect. A larger number makes non-randomisation less of a concern.
@@ B.5.2 Number of points
B.5.3 Impact of treatment rate